# Geolocation with Real Human Gameplay Data: A Large-Scale Dataset and Human-Like Reasoning Framework

### v2

@@ 3.3. Human Player Performance
4. Geographic Chain of Thought
In this section, we introduce GeoCoT, a novel chain-of-thought prompting framework for graph-based and geolocation tasks. Unlike standard CoT prompting which performs generic step-by-step reasoning, GeoCoT introduces a domain-specific, hierarchically structured reasoning process that mimics how humans localize geographic information from broad regions to fine-grained details.
4.1. Rethinking Geolocation Task
As discussed in § 2.1, the geolocation task has traditionally relied on classification-based (Clark et al., 2023b; Pramanick et al., 2022; Weyand et al., 2016b) and retrieval-based methods (Zhu et al., 2022; Zhang et al., 2023b), as shown in Figure 3. While these approaches have advanced the field, they face significant limitations in precision and scalability, prompting a rethinking of the task.
Inspired by how humans gradually narrow down locations from broad to fine-grained observations (Luo et al., 2022), we propose a new geolocation paradigm: predicting geographic locations through a step-by-step reasoning process. Unlike traditional approaches limited by grid-based classification and exhaustive databases, our model generates natural language reasoning, guiding it to the final predicted city. To implement this paradigm, we introduce GeoCoT (Geographic Chain-of-Thought), a framework designed for both interpretability and accuracy.
4.2. GeoCoT Deisgn
@@
It is important to note that GeoCoT does not require any concrete knowledge about the specific features of locations. Instead, it offers reasoning tutorials designed to help LVMs identify geographic clues by leveraging their existing knowledge.
5. Experiments
In this section, we first introduce our experimental settings, then evaluate GeoCoT in terms of its general geolocalization ability, followed by a detailed evaluation of its reasoning process.
5.1. Setting
@@
We selected 500 geo-tagged locations with high inferential value from the dataset to serve as a test set, using a stratified sampling method across continents to ensure balanced geographic distribution. This number is larger than in previous works (Liu et al., 2023b; Guan et al., 2024), which typically include only a few dozen case studies to examine the reasoning process. We define ”high inferential value” as locations with moderate difficulty—challenging enough to be correctly identified by experienced participants, but not so easy that beginners can do so effortlessly. Specifically, we selected 20 mainstream countries across six continents as representative samples and extracted tasks with an average player score of around 3,000 out of 5,000 for annotation. This test set has been publicly released on GitHub. Our GeoCoT framework is implemented using GPT-4o.
5.2. Baselines
We compare our model, GeoCoT, against several strong baselines representing the latest advancements in geolocation on our dataset. General Open-Source VLMs: LLaVA-1.6 (Liu et al., 2023a) utilizes a fully connected vision-language connector, effectively bridging visual inputs with linguistic features to deliver strong results in geolocation tasks. Llama-3.2-vision (Meta, 2024) demonstrates advanced multi-modal reasoning capabilities, making it a powerful open-source vision-language model. Qwen-VL (Bai et al., 2023), leveraging vast datasets of billions of image-text pairs, achieves robust performance in geolocation through its strong visual and spatial semantic understanding. Baselines Targeting Geolocation Tasks: GeoCLIP (Vivanco Cepeda et al., 2024), inspired by CLIP, aligns images with GPS coordinates using a retrieval-based approach to enhance geolocation. GeoReasoner (Li et al., 2024a) combines geospatial reasoning with visual-language alignment for state-of-the-art geolocation performance. Closed-Source VLMs: GPT-4o (OpenAI, 2024) excels in vision reasoning tasks with its advanced multi-modal capabilities. Furthermore, GPT-4o(CoT), following the setting of cot-zero-shot (Wei et al., 2022), leverages chain-of-thought reasoning to improve performance in complex scenarios. All models are evaluated using the same input format and test set to ensure a fair comparison.
5.3. Overall Performance Evaluation of GeoCoT
@@
In this subsection, we evaluate the city location prediction performance of our model in comparison with the latest LVM models. We evaluate geolocation performance from two aspects: first, location prediction compared with the ground truth at various levels; and second, the direct calculation of the Earth’s surface distance. We present the location prediction performance in Table 2, evaluated across three levels: city, country, and continent. Performance is measured using accuracy, which calculates the proportion of correct predictions out of all predictions; recall, which determines the proportion of true positive predictions out of all actual positive cases; and the F1 score, which balances precision and recall to provide their harmonic mean.
@@
The results reveal several key observations. First, open-source LVMs such as LLaMA-3.2-Vision achieve competitive performance, performing on par with GPT-4o and GPT-4o (CoT), demonstrating their effectiveness in location prediction tasks. Second, performance varies across different levels of granularity. While GPT-4o (CoT) ranks second at the city level, it underperforms at the country level, highlighting the importance of multi-level evaluation to fully assess a model’s geolocation reasoning ability. Finally, our model, GeoCoT, consistently achieves top performance across all nine metrics and three levels, demonstrating its robustness and adaptability in geolocation tasks. Additionally, GeoCLIP surpasses GPT-4o at the continent level, which can be attributed to its pretraining on image-GPS pairs, making it particularly well-suited for coarse-grained geolocation tasks. Coarse-grained continent-level predictions typically require less detailed local knowledge and instead rely on broader geographic cues, such as climate, landscapes, and cultural markers. However, GeoCLIP performs poorly at finer granularities like country and city levels, suggesting that it lacks a strong capability for geographic reasoning beyond direct visual features.
Table 3. Accuracy of different models on geolocation tasks at various scales. Numbers in bold mean that the improvement to the best baseline is statistically significant (a two-tailed paired t-test with p-value <0.01).
@@
Next, in Table 3, we present the accuracy of each model by measuring the geographic distance between the predicted city and the ground truth. The metrics represent the proportion of predictions within three distance thresholds: Street (1 km), City (25 km), and Country (750 km). Higher values indicate better performance, with stricter thresholds assessing fine-grained localization and larger thresholds evaluating coarse-level accuracy. The results show that GPT-4o and Llama-3.2-vision outperform the dedicated large-scale model GeoCLIP for geolocation, even under finer-grained evaluation settings. For example, at the street-level threshold, GPT-4o achieves 0.045 compared to GeoCLIP’s 0.035, and at the city-level threshold, GPT-4o scores 0.147, nearly double GeoCLIP’s 0.077. Moreover, our proposed GeoCoT paradigm demonstrates even greater improvements. At the street level, GeoCoT achieves 0.073, significantly outperforming both GeoCLIP (0.035) and GPT-4o (0.045). Similarly, at the city level, GeoCoT achieves 0.157, and at the country level (750 km), it achieves 0.711, the highest among all models. These results highlight GeoCoT’s strong performance and the potential of its reasoning framework for geolocation tasks.
5.4. GeoEval: Reference-Based Evaluation of GeoCoT Reasoning
Beyond evaluating overall task performance, we focus on analyzing the reasoning process of GeoCoT, which emulates a human-like reasoning approach. To establish a reference for this evaluation, three gaming enthusiasts collaboratively constructed reasoning processes for the same 500 cases based on geo-tagged locations. We designated these as the reasoning ground truth (a human-annotated example can be found in Appendix C). These GT annotations serve as a benchmark within our evaluation framework, GeoEval. The evaluation process utilizes (1) GPT-based assessment through GPTScore (Fu et al., 2023) and (2) prompt-based scoring.
Our prompt-based scoring includes four dimensions ranging from 0-5, and the detailed prompts can be found in Github. The first dimension is the completeness of feature extraction (CE), which evaluates whether all key clues provided in the GT are comprehensively covered and accurately described in the reasoning process. Comprehensive feature extraction ensures that reasoning outcomes are based on sufficient factual evidence, thereby enhancing their reliability and accuracy. The second dimension is the accuracy of feature extraction (AE), which measures whether the identified and described attributes or characteristics of the key information in the GT are correct. Misidentified features can lead to reasoning outcomes that deviate from the facts, reducing the credibility of the results. The third dimension is the accuracy of reasoning and cue correspondence (AC), which assesses whether the reasoning process derives reasonable conclusions based on the extracted cues and maintains consistency with the reasoning logic presented in the GT. Incorrect correspondence between cues and conclusions can result in outcomes that deviate from reality. The final dimension is the logical coherence of reasoning (LC), which evaluates the consistency, logical flow, and adherence to common sense within the reasoning chain. Logical errors compromise the reliability of the reasoning process and hinder the model’s ability to arrive at accurate conclusions.
@@ 5.4. GeoEval: Reference-Based Evaluation of GeoCoT Reasoning
Table 4. Evaluation of GeoCoT’s reasoning process using ground truth-based metrics within the GeoEval framework. Numbers in bold mean that the improvement to the best baseline is statistically significant (a two-tailed paired t-test with p-value <0.01).
@@
The experimental results in Table 4 highlight the significant advantages of GeoCoT compared to baseline models across all evaluation metrics. GeoCoT achieves the highest GPTScore of 0.728, outperforming GPT-4o (CoT) (0.663) and -1.6 (0.478), demonstrating its superior alignment with human-constructed reasoning processes. In terms of feature extraction, GeoCoT achieves a CE score of 2.690 and an AE score of 3.538, significantly surpassing GPT-4o (CoT) and the dedicated GeoReasoner model. Furthermore, GeoCoT’s performance in reasoning accuracy and logical coherence is unmatched, with AC and LC scores of 3.696 and 3.945, compared to GPT-4o (CoT), which scores 3.156 and 3.540, and GeoReasoner, which lags behind at 1.719 and 2.038. These results clearly demonstrate that GeoCoT not only captures key information more comprehensively but also maintains a more accurate and logically coherent reasoning process compared to both reasoning-based models and traditional baselines like GeoReasoner.
5.5. Intrinsic Evaluation of GeoCoT Reasoning
Figure 4. Qualitative comparison of LLaVA, GPT4o, and GeoReasoner. Clues are shown in blue, correct predictions in green, incorrect in red, and vague/uncertain guesses in orange.
We begin with a ground truth-based evaluation, comparing GeoCoT’s reasoning to human-authored processes to assess its alignment with established reasoning patterns. To complement this, we conduct an intrinsic evaluation focused on hallucination errors—assessing logical consistency, coherence, and robustness without relying on external references. Given the need for multimodal judgment, this evaluation is performed by human annotators.
Following previous work on assessing hallucinations in terms of objects, attributes, and relationships (Li et al., 2023a; Sun et al., 2023), we evaluate the quality of synthetic data across three key dimensions: (1) Object Hallucination (OH): assesses whether the synthetic data includes objects that do not exist in the image. Object Hallucination evaluates the extent to which synthetic data introduces fictional elements. (2) Fact Hallucination (FH) measures the accuracy of factual information within the synthetic data. Fact Hallucination occurs when the synthetic data contains facts, figures, or other information that is incorrect or not supported by the original data. (3) Attribution Hallucination (AH) evaluates whether the synthetic data incorrectly attributes properties, characteristics, or relations to entities or objects. To quantify hallucinations, each detected error is counted as one instance in the corresponding dimension. To evaluate these dimensions, we invited 2 human annotators with professional backgrounds in geographic reasoning and data validation to assess GPT-4o, GeoReasoner, and our proposed GeoCoT model. These three baselines provide textual reasoning processes across 1,500 evaluated cases. The results, shown in Table 5, indicate the number of errors in each dimension, demonstrating that GeoCoT significantly reduces hallucination errors compared to the other models. The inter-annotator agreement, measured by Cohen’s Kappa, is 0.82 for OH, 0.79 for FH, and 0.85 for AH, indicating substantial agreement across all dimensions. The correlation between the human evaluation scores and the automatic evaluation metrics in Table 2 is –0.99 (p ¡ 0.01), demonstrating a strong inverse relationship: as hallucination errors decrease, overall geolocation performance improves significantly.
Table 5. Hallucination Evaluation on Reasoning Data.
@@ 5.5. Intrinsic Evaluation of GeoCoT Reasoning
Table 6. Performance comparison of GeoCoT and state-of-the-art geolocation models on traditional benchmarks. Numbers in bold mean that the improvement to the best baseline is statistically significant (a two-tailed paired t-test with p-value <0.01).
Model Im2GPS Im2GPS3K
Street City Country Street City Country
1km 25km 750km 1km 25km 750km
LLaVA-1.6 0.04 0.18 0.39 0.03 0.14 0.32
Llama-3.2-Vision 0.09 0.37 0.65 0.07 0.27 0.52
Qwen-VL 0.04 0.21 0.37 0.04 0.15 0.26
GeoCLIP 0.17 0.41 0.77 0.13 0.32 0.67
GeoReasoner 0.05 0.19 0.33 0.04 0.15 0.26
GPT-4o 0.13 0.47 0.74 0.14 0.40 0.66
GPT-4o(CoT) 0.16 0.49 0.77 0.14 0.45 0.69
GeoCoT 0.22 0.55 0.83 0.15 0.46 0.74
5.6. Case Study
We present two examples in Figure 4 to analyze the performance of LLaVA, GPT4o, and GeoReasoner, highlighting the effectiveness of our GeoCoT approach. In the first example, struggles to provide a specific prediction, reflecting its reliance on general architectural cues and its tendency to consider broad regions such as the United Kingdom or France. GPT4o, despite identifying key features of the European landscape, incorrectly associates them with Germany, indicating limitations in handling specific regional markers. In contrast, GeoCoT accurately pinpoints the location in France by effectively integrating textual clues, architectural elements, and environmental context.
In the second example, GeoCoT correctly identifies the location as San Francisco, USA, by analyzing U.S. traffic standards, license plates, and local signage, demonstrating strong contextual reasoning. LLaVA-1.6 makes a broad prediction, covering the U.S., Australia, and the U.K., showing uncertainty from general cues. GPT-4o misidentifies the scene as Seattle, relying on architectural similarities but missing key details.
5.7. Generalizability Evaluation
Even though our dataset is more comprehensive and human-annotated, we are also interested in evaluating how our model performs on traditional geolocation datasets to provide a more thorough comparison. Hence, we select two existing benchmark datasets, Im2GPS (Hays and Efros, 2008) and Im2GPS3K (Vo et al., 2017b), due to their popularity and widespread use in geolocation tasks as standard benchmarks for evaluating model performance. Similarly, we use the center point coordinates of the city text address in GeoCoT’s output and measure the distance between the output and the ground truth locations.
We present the performance results in Table 6. We observe that state-of-the-art geolocation models, such as GeoCLIP, perform well on traditional geolocation tasks, surpassing GPT-4o and coming close to our model, GeoCoT. However, this is in contrast to the results shown in Table 2, where GeoCLIP significantly underperforms GPT-4o on fine-grained city- and country-level geolocation tasks. This discrepancy suggests that these baseline models may be overfitting to the specific datasets they were trained on, lacking the generalization ability required for more diverse or fine-grained geolocation challenges. In contrast, our model consistently outperforms traditional methods across different granularity levels and datasets without any training, and thus does not suffer from overfitting.
6. Conclusion
In this work, we present the largest geolocation dataset to date, collected from a geolocation game platform with 740K users over two years. The dataset comprises 25M entries of metadata, including 3M geo-tagged locations spanning most of the globe, each annotated thousands to tens of thousands of times by human users. This dataset enables diverse difficulty-level analysis and highlights the limitations of current LVMs. We also introduce a generation-based reasoning solution for the geolocation task, where the LVM generates reasoning chains by leveraging clues from images and produces the final predicted location. Using our GeoEval set of metrics, we demonstrate that our GeoCoT framework significantly outperforms state-of-the-art general and task-specific baselines on this dataset.
In future work, we plan to enhance model interpretability and robustness, explore multi-modal integration of text and visuals, and expand the dataset to better cover underrepresented regions for improved global coverage and fairness.

### v3

@@ 3.3 Human Annotations
Kimi-latest 0.027 0.023 0.024 0.135 0.112 0.156 0.315 0.148 0.220
Kimi-latest(CoT) 0.028 0.029 0.029 0.139 0.092 0.124 0.331 0.154 0.210
Kimi-latest(GeoCoT) 0.038 0.035 0.029 0.239 0.192 0.224 0.361 0.168 0.232
@@
Table 2: Comparison of Precision, Recall and F1 scores in country-level and city-level geolocation. The scores are represented as follows: best, second, and third. Numbers in bold mean that the improvement to the best baseline is statistically significant (a two-tailed paired t-test with p-value <0.05).
4 Geographic Chain of Thought
4.1 Rethinking Geolocation Task
As discussed in § 2.1, the geolocation task has traditionally relied on classification-based Clark et al. (2023b); Pramanick et al. (2022); Weyand et al. (2016b) and retrieval-based methods Zhu et al. (2022); Zhang et al. (2023a), as shown in Figure 2. While these approaches have advanced the field, they face significant limitations in precision and scalability, prompting a rethinking of the task. Inspired by how humans gradually narrow down locations from broad to fine-grained observations Luo et al. (2022), we propose a new geolocation paradigm: predicting geographic locations through a step-by-step reasoning process. Unlike traditional approaches limited by grid-based classification and exhaustive databases, our model generates natural language reasoning, guiding it to the final predicted city. To implement this paradigm, we introduce GeoCoT (Geographic Chain-of-Thought), a framework designed for both interpretability and accuracy.
4.2 GeoCoT Deisgn
@@
It is important to note that GeoCoT does not require any concrete knowledge about the specific features of locations. Instead, it offers reasoning tutorials designed to help LVMs identify geographic clues by leveraging their existing knowledge. Our inference time and efficiency analysis in Appendix E shows that GeoCoT incurs only about a 10% increase in inference latency compared to standard CoT. We further provide an ablation study in Appendix F.
5 Experiments
5.1 Setting
@@
We selected 500 geo-tagged locations with high inferential value from the dataset to serve as a test set, using a stratified sampling method across continents to ensure balanced geographic distribution. This number is larger than in previous works Liu et al. (2023b); Guan et al. (2024), which typically include only a few dozen case studies to examine the reasoning process. Specifically, we selected 20 mainstream countries across six continents as representative samples and extracted tasks with an average player score of around 3,000 out of 5,000 for annotation. This test set has been publicly released on GitHub.
5.2 Baselines
We compare GeoCoT against strong baselines representing recent geolocation advancements. For general open-source VLMs, we evaluate LLaVA-1.6 Liu et al. (2023a), Llama-3.2-vision Meta (2024), and Qwen2.5-VL Bai et al. (2023), also include geolocation-specific models GeoCLIP Vivanco Cepeda et al. (2024) and GeoReasoner Li et al. (2024). Furthermore, we assess closed-source performance using GPT-4o OpenAI (2024), Kimi k2  Team et al. (2025) and also utilizing chain-of-thought reasoning Wei et al. (2022). All models are evaluated using the same input format and test set to ensure a fair comparison.
5.3 Overall Performance Evaluation of GeoCoT
@@
We evaluate geolocation performance from two aspects: first, location prediction compared with the ground truth at various levels; and second, the direct calculation of the Earth’s surface distance. We present the location prediction performance in Table 2, evaluated across three levels: city, country, and continent. Performance is measured using accuracy, which calculates the proportion of correct predictions out of all predictions; recall, which determines the proportion of true positive predictions out of all actual positive cases; and the F1 score, which balances precision and recall to provide their harmonic mean.
@@
The results reveal several key observations. First, open-source LVMs such as LLaMA-3.2-Vision achieve competitive performance, performing on par with GPT-4o and GPT-4o (CoT), demonstrating their effectiveness in location prediction tasks. Second, performance varies across different levels of granularity. While GPT-4o (CoT) ranks second at the city level, it underperforms at the country level, highlighting the importance of multi-level evaluation to fully assess a model’s geolocation reasoning ability. Finally, our model, GeoCoT, consistently achieves top performance across all nine metrics and three levels, demonstrating its robustness and adaptability in geolocation tasks. Additionally, GeoCLIP surpasses GPT-4o at the continent level, which can be attributed to its pretraining on image-GPS pairs, making it particularly well-suited for coarse-grained geolocation tasks. Coarse-grained continent-level predictions typically require less detailed local knowledge and instead rely on broader geographic cues, such as climate, landscapes, and cultural markers. However, GeoCLIP performs poorly at finer granularities like city levels, suggesting that it lacks a strong capability for geographic reasoning beyond direct visual features.
@@
Table 3: Accuracy of different models on geolocation tasks at various scales. Numbers in bold mean that the improvement to the best baseline is statistically significant (a two-tailed paired t-test with p-value <0.05).
Next, in Table 3, we present the accuracy of each model by measuring the geographic distance between the predicted city and the ground truth. The metrics represent the proportion of predictions within three distance thresholds: Street (1 km), City (25 km), and Country (750 km). Higher values indicate better performance, with stricter thresholds assessing fine-grained localization and larger thresholds evaluating coarse-level accuracy. The results show that GPT-4o and Llama-3.2-vision outperform the dedicated large-scale model GeoCLIP for geolocation, even under finer-grained evaluation settings. For example, at the street-level threshold, GPT-4o achieves 0.045 compared to GeoCLIP’s 0.035, and at the city-level threshold, GPT-4o scores 0.147, nearly double GeoCLIP’s 0.077. Moreover, our proposed GeoCoT paradigm demonstrates even greater improvements. At the street level, GeoCoT achieves 0.073, significantly outperforming both GeoCLIP (0.035) and GPT-4o (0.045). Similarly, at the city level, GeoCoT achieves 0.157, and at the country level (750 km), it achieves 0.711, the highest among all models. These results highlight GeoCoT’s strong performance and the potential of its reasoning framework for geolocation tasks. We further provide an ablation study on GeoCoT in Appendix F.
Model Im2GPS Im2GPS3K
Street City Country Street City Country
1km 25km 750km 1km 25km 750km
LLaVA-1.6 0.04 0.18 0.39 0.03 0.14 0.32
Llama-3.2-Vision 0.09 0.37 0.65 0.07 0.27 0.52
Qwen-VL 0.04 0.21 0.37 0.04 0.15 0.26
GeoCLIP 0.17 0.41 0.77 0.13 0.32 0.67
GeoReasoner 0.05 0.19 0.33 0.04 0.15 0.26
PlaNet 0.08 0.25 0.54 0.09 0.25 0.48
GPT-4o 0.13 0.47 0.74 0.14 0.40 0.66
GPT-4o(CoT) 0.16 0.49 0.77 0.14 0.45 0.69
GeoCoT 0.22 0.55 0.83 0.15 0.46 0.74
Table 4: Performance comparison of GeoCoT and state-of-the-art geolocation models on traditional benchmarks. Numbers in bold mean that the improvement to the best baseline is statistically significant (a two-tailed paired t-test with p-value <0.05).
5.4 Generalizability Evaluation
Even though our dataset is more comprehensive and human-annotated, we are also interested in evaluating how our model performs on traditional geolocation datasets to provide a more thorough comparison. Hence, we select two existing benchmark datasets, Im2GPS Hays and Efros (2008) and Im2GPS3K Vo et al. (2017b), due to their popularity and widespread use in geolocation tasks as standard benchmarks for evaluating model performance. We present the performance results in Table 4. We observe that state-of-the-art geolocation models, such as GeoCLIP, perform well on traditional geolocation tasks, surpassing GPT-4o and coming close to our model, GeoCoT. However, this is in contrast to the results shown in Table 2, where GeoCLIP significantly underperforms GPT-4o on fine-grained city- and country-level geolocation tasks. This discrepancy suggests that these baseline models may be overfitting to the specific datasets they were trained on, lacking the generalization ability required for more diverse or fine-grained geolocation challenges.
5.5 GeoEval: Reference-Based Evaluation
Beyond evaluating overall task performance, we focus on analyzing the reasoning process of GeoCoT, which emulates a human-like reasoning approach. To establish a reference for this evaluation, three gaming enthusiasts collaboratively constructed reasoning processes for the same 500 cases based on geo-tagged locations. We designated these as the reasoning ground truth (a human-annotated example can be found in Appendix C). These GT annotations serve as a benchmark within our evaluation framework, GeoEval. The evaluation process utilizes GPT-based assessment through GPTScore Fu et al. (2023) and prompt-based scoring.
Our prompt-based scoring includes four dimensions (scored 0-5; detailed prompts are on Github). First, completeness of feature extraction (CE) evaluates whether all key clues in the GT are comprehensively covered, ensuring reasoning relies on sufficient factual evidence. Second, accuracy of feature extraction (AE) measures the correctness of identified attributes to prevent reasoning deviations caused by misidentified features. Third, accuracy of reasoning and cue correspondence (AC) assesses whether conclusions derived from extracted cues are reasonable and consistent with the GT logic. Finally, logical coherence of reasoning (LC) evaluates the consistency, flow, and adherence to common sense within the reasoning chain, ensuring the reliability of the conclusion.
@@ 5.5 GeoEval: Reference-Based Evaluation
Table 5: Evaluation of GeoCoT’s reasoning process using ground truth-based metrics within the GeoEval framework. Numbers in bold mean that the improvement to the best baseline is statistically significant (a two-tailed paired t-test with p-value <0.05).
The experimental results in Table 5 highlight the significant advantages of GeoCoT compared to baseline models across all evaluation metrics. GeoCoT achieves the highest GPTScore of 0.728, outperforming GPT-4o (CoT) (0.663) and LLaVA-1.6 (0.478), demonstrating its superior alignment with human-constructed reasoning processes. In terms of feature extraction, GeoCoT achieves a CE score of 2.690 and an AE score of 3.538, significantly surpassing GPT-4o (CoT) and the dedicated GeoReasoner model. These results clearly demonstrate that GeoCoT not only captures key information more comprehensively but also maintains a more accurate and logically coherent reasoning process compared to both reasoning-based models and traditional baselines like GeoReasoner.
5.6 Intrinsic Evaluation of GeoCoT Reasoning
We begin with a ground truth-based evaluation, comparing GeoCoT’s reasoning to human-authored processes to assess its alignment with established reasoning patterns. To complement this, we conduct an intrinsic evaluation focused on hallucination errors, coherence, and robustness without relying on external references. Given the need for multimodal judgment, this evaluation is performed by human annotators.
@@ 5.6 Intrinsic Evaluation of GeoCoT Reasoning
Table 6: Hallucination Evaluation on Reasoning Data.
Following previous work on assessing hallucinations in terms of objects, attributes, and relationships Li et al. (2023); Sun et al. (2023), we evaluate the quality of synthetic data across three key dimensions: (1) Object Hallucination (OH): assesses whether the synthetic data includes objects that do not exist in the image. Object Hallucination evaluates the extent to which synthetic data introduces fictional elements. (2) Fact Hallucination (FH) measures the accuracy of factual information within the synthetic data. Fact Hallucination occurs when the synthetic data contains facts, figures, or other information that is incorrect or not supported by the original data. (3) Attribution Hallucination (AH) evaluates whether the synthetic data incorrectly attributes properties, characteristics, or relations to entities or objects. To quantify hallucinations, each detected error is counted as one instance in the corresponding dimension. To evaluate these dimensions, we invited 2 human annotators with professional backgrounds in geographic reasoning and data validation to assess GPT-4o, GeoReasoner, and our proposed GeoCoT model. These three baselines provide textual reasoning processes across 1,500 evaluated cases. The results, shown in Table 6, indicate the number of errors in each dimension, demonstrating that GeoCoT significantly reduces hallucination errors compared to the other models. The inter-annotator agreement, measured by Cohen’s Kappa, is 0.82 for OH, 0.79 for FH, and 0.85 for AH, indicating substantial agreement across all dimensions. The detailed reasoning case studies are listed in Appendix D.
6 Conclusion
In this work, we advance the field of image geolocation by introducing GeoComp, the largest human-annotated dataset to date, comprising 2.7 million locations and 25 million real-world gameplay records. Leveraging this rich data, we propose GeoCoT, a geographical chain-of-thought framework that mimics human reasoning to enhance LVM capabilities. Comprehensive evaluations using our novel GeoEval metric demonstrate that GeoCoT significantly outperforms state-of-the-art baselines in both accuracy and interpretability.
Limitations
Despite the advancements presented, our work has three main limitations. First, although GeoComp is globally diverse, geographic coverage remains uneven, with sparse representation in regions like Africa and parts of Oceania due to limited map service availability. Nevertheless, compared to existing large-scale geolocation benchmarks (e.g., OSV-5M), GeoComp substantially mitigates extreme regional bias and achieves a markedly more balanced global distribution, as demonstrated by the Geo-Pop Index analysis. Second, GeoCoT relies on computationally intensive LVM (e.g., GPT-4o). The multi-step reasoning process increases inference latency and cost compared to lightweight, embedding-based retrieval methods, potentially limiting real-time deployment. While this overhead is inherent to CoT-style reasoning, our empirical analysis (Appendix E) indicates that GeoCoT does not incur substantially higher latency than standard CoT prompting across different models. Finally, while GeoCoT excels at city-level localization through reasoning, it may struggle with precise meter-level coordinate prediction in non-urban environments where semantic cues (text, landmarks) are absent, an area where retrieval-based methods still hold an advantage. This limitation stems from the intrinsic reliance of reasoning-based methods on interpretable visual cues; in such low-semantic settings, direct visual matching remains more effective, and we view the integration of reasoning-based and retrieval-based paradigms as a promising direction for future work.